ABC Company offers two products. The following represents the results of a month's operations:

Product A: Revenues $100,000. Total Variable Costs $50,000. Units sold 100,000

Product B: Revenues $100,000. Total Variable Costs $60,000. Units sold 50,000

Total Fixed Costs          $60,000

1. Find the break-even point in units of A and B.
Answer #1

Product A:

FC = 60,000

Sales price per unit = Revenues / Units sold = 100000 / 100,000 = 1

Variable costs per unit = Total Variable Costs / No. of units = 50,000 / 100,000 = 0.50

Break-Even point (units) = Fixed Costs / (Sales price per unit – Variable costs per unit)

Break-Even point (units) = 60,000 / (1 - 0.50) = 120,000 units

 

Product B:

FC = 60,000

Sales price per unit = Revenues / Units sold = 100000 / 100,000 = 1

Variable costs per unit = Total Variable Costs / No. of units = 60,000 / 100,000 = 0.60

Break-Even point (units) = Fixed Costs / (Sales price per unit – Variable costs per unit)

Break-Even point (units) = 60,000 / (1 - 0.60) = 150,000 units